OpenStudy (anonymous):

Solve for x: −3(x + 3) = −3(x + 1) − 5 6 −6 All real numbers No solution

OpenStudy (anonymous):

-3x - 9 = -3x - 3 - 5 -3x + 3x = -3 - 5 - 9 => 0 = -17 Weird sum eh?

OpenStudy (anonymous):

Yes, So No Solution

OpenStudy (anonymous):

I think the answer should be no solution because such an equation is not possible.